# BitNet v2: Native 4-bit Activations with Hadamard Transformation for 1-bit LLMs

### v1

@@ 3.3 Ablation Study
No rotatation W1.58A8 diverged diverged
Weight & activation rotatation 50.47 11.14 55.55 9.69
Activation rotatation 51.16 11.14 55.71 9.72
No rotatation W1.58A4 diverged diverged
Weight & activation rotatation 50.09 11.33 54.98 9.81
Activation rotatation 50.41 11.33 55.43 9.85
@@ Appendix A More results
No rotatation 1.3B diverged
Weight & activation rotatation 27.13 50.29 48.32 69.04 54.30 53.76 50.47
Activation rotatation 27.90 49.96 48.37 69.42 57.22 54.14 51.17
No rotatation 3B diverged
Weight & activation rotatation 30.03 55.72 56.81 71.65 59.43 59.65 55.54
Activation rotatation 30.55 55.56 57.19 71.33 58.72 60.90 55.71
@@
No rotatation 1.3B diverged
Weight & activation rotatation 27.22 49.12 47.77 69.37 54.54 52.49 50.09
Activation rotatation 27.56 49.58 48.00 68.23 55.49 53.58 50.41
No rotatation 3B diverged
Weight & activation rotatation 29.44 54.46 56.57 71.93 57.85 59.64 54.98
Activation rotatation 28.92 55.01 56.59 71.65 59.67 60.74 55.43

### v2

@@ 3.3 Ablation Study
No rotation W1.58A8 51.01 11.02 55.22 9.71
Weight & activation rotation 50.47 11.14 55.55 9.69
Activation rotation 51.16 11.14 55.71 9.72
No rotation W1.58A4 diverged diverged
Weight & activation rotation 50.09 11.33 54.98 9.81
Activation rotation 50.41 11.33 55.43 9.85
@@ Appendix A More results
No rotation 1.3B diverged
Weight & activation rotation 27.13 50.29 48.32 69.04 54.30 53.76 50.47
Activation rotation 27.90 49.96 48.37 69.42 57.22 54.14 51.17
No rotation 3B diverged
Weight & activation rotation 30.03 55.72 56.81 71.65 59.43 59.65 55.54
Activation rotation 30.55 55.56 57.19 71.33 58.72 60.90 55.71
@@
No rotation 1.3B diverged
Weight & activation rotation 27.22 49.12 47.77 69.37 54.54 52.49 50.09
Activation rotation 27.56 49.58 48.00 68.23 55.49 53.58 50.41
No rotation 3B diverged
Weight & activation rotation 29.44 54.46 56.57 71.93 57.85 59.64 54.98
Activation rotation 28.92 55.01 56.59 71.65 59.67 60.74 55.43